In triangle ABC, AB is 9 cm. shorter that AC, while BC is 3 cm. longer than AC. If the perimeter of the triangle is 48 cm. find the lengths of the three sides.
Answer by Annabelle1(69)

let the length of AC be X
so.....
AC=x
BC=x+3
AB=x-9
adding the sides together to find the perimeter
x+ (x+3) + (x-9) = 48
3x-6=48
3x=54
x=18
so AC=18
BC= 18+3=21
AB=18-9=9